Suppose a firm faces a total cost function of: T C(Q) = 55 − 2Q + 10Q 2 (a) How would you express the average fixed cost for any level of output? What is the average fixed cost for 5 units of output? (b) How would you express the variable costs? What are variable costs when output is 6? (c) What is the marginal cost for the 12th unit of output? (d) At what level of output do marginal costs begin to exceed average total costs? If marginal costs never exceed average total costs, make that clear.
Answer #1

TCCO )= 55-20 +100² 6) TFC - 55 AFC-55 [:AFC - T&T When Q = 5, AFC = 38 = 11 (6) TVC= – 2Q + 10Q² When Q=6, TVC - 206 ) +10 (